OpenStudy (anonymous):

A system consist of N no of particles ,N>>1.each particle can have only one of the two energies E1or E2+ε(ε>0).if the system is in equllibrium at a temperature T,the avg no. of particles with energy E1 is.....

OpenStudy (michele_laino):

the requested probability, is given by the subsequent formula: \[\huge p = \frac{{{e^{ - {E_1}/kT}}}}{{{e^{ - {E_1}/kT}} + {e^{ - \left( {{E_1} + \varepsilon } \right)/kT}}}} = ...?\] where \(E_1\) and \(E_1+\epsilon\), are the energies of such quantum states

OpenStudy (michele_laino):

it is the probability that the system lies in quantum state with energy \(E_1\)

OpenStudy (michele_laino):

with such probability, it is easy to compute the requested average number of particles. Furthermore, \(k\) is the \(Boltzmann\) constant

OpenStudy (anonymous):

sry bt i didnt get this .....wht is d ans & will u tell plz me wht does question represents.......

OpenStudy (anonymous):

options are like....1.N/2 2.N/e^(ε/KT+1) 3.N/e^(-ε/KT+1) 4.Ne^(-ε/KT)

OpenStudy (michele_laino):

the requested number is given by the subsequent product: \[\Large \begin{gathered} Np = \frac{{N{e^{ - {E_1}/kT}}}}{{{e^{ - {E_1}/kT}} + {e^{ - \left( {{E_1} + \varepsilon } \right)/kT}}}} = \hfill \\ \hfill \\ = N{e^{ - {E_1}/kT}}\frac{1}{{{e^{ - {E_1}/kT}}\left( {1 + {e^{ - \varepsilon /kT}}} \right)}} = ...? \hfill \\ \end{gathered} \] please continue @pratima99